For each of the thin lenses (L1 and L2 ) shown in the fig
Answer #1


1/f_1a = (n-1)[(1/R1)+(1/R2)] = 0.0867
f1a = 11.53 cm
f1b = 11.53 cm..

B) 1/f1b = (1-n)/R2 = -0.5/8.6 = -17.2 cm...

f2b = 17.2 cm..

C) 1/f_1c = (n-1)[(1/R1)-(1/R2)] = -0.00925 ...
f_1c = -108.1 cm...
f2_c = 108.1 cm...

D) 1/f1d = (1-n)/R1 = -0.5/9.15..

f1d = -18.3 cm...

f2d = 18.3cm...

E) 1/f_1e = (1-n)[(1/R1)+(1/R2)] = -0.08977...
f1e = -11.13 cm...

f2e = -11.13 cm